# Multi-platoon car-following models with flexible platoon sizes and communication levels

### v2

@@ Abstract
sIn this paper, we extend a single platoon car-following (CF) model to some multi-platoon CF models for connected and autonomous vehicles (CAVs) with flexible platoon size and communication level. Specifically, we consider forward and backward communication methods between platoons with delays. Some general results of linear stability are mathematically proven, and numerical simulations are performed to illustrate the effects of platoon sizes and communication levels, as well as to demonstrate the potential for stabilizing human-driven vehicles (HDVs) in mixed traffic conditions. The simulation results are consistent with theoretical analysis, and demonstrate that in the ring road scenario, CAV platoons can stabilize certain percentage of HDVs. This paper can provide suggestions for the design of communication system of autonomous vehicles (AVs), and management of mixed traffic flow of CAVs and HDVs.
Index Terms:
CAV platoon, vehicular communication, linear stability, numerical simulation
I Introduction
@@
The next stage of CF models includes multi-vehicle interaction, moving beyond the classic models considering only the immediate leader, thus modeling traffic flow more realistically.CF patterns such as multi-following [9] and backward-following [10], both based on OVM, have been increasingly explored and have inspired subsequent research on mixed traffic flow [11], [12], where CAVs are designed to better stabilize traffic. In fully autonomous traffic scenarios, various platooning designs can be integrated to enhance both efficiency and safety [13], [14], [15], [16, 17], [18]. Further improvements to CF models include the addition of delay factors [19], [20], as well as the integration of control mechanisms such as safety-prioritized control [21, 22] and feedback control [23]. These extended CF models, coupled with advanced control strategies, reflect the growing trend on improving traffic stability, efficiency, and safety. By accounting for multi-vehicle interactions and integrating autonomous systems, these models contribute significantly to enhancing traffic performance, particularly in mixed autonomy scenarios.
@@
Control strategies for groups of CAVs, particularly ACC and Cooperative Adaptive Cruise Control (CACC), have been extensively studied to improve vehicle platooning efficiency and safety [24, 25]. ACC primarily focuses on maintaining safe distances between vehicles by adjusting speed based on sensor data, achieving better results than typical human drivers. However, it operates in a decentralized manner without relying on vehicle-to-vehicle (V2V) communication. On the other hand, CACC utilizes V2V communication to enable more precise control and coordination among CAV platoons [26, 27, 28]. Additionally, the quality of communications plays a crucial role in the stability of CAV platoons, as shown in studies on robust communication and stability analysis [29, 30, 31]. These studies highlight the advantages of CACC over ACC, particularly regarding communication and coordination within platoons.
@@
Beyond ACC and CACC, advanced control strategies such as Model Predictive Control (MPC), reinforcement learning (RL), and stochastic optimization have been explored to further enhance the performance of CAV platoons in complex traffic scenarios. MPC has been widely used due to its ability to predict future states and optimize control actions over a finite horizon [32, 33, 34]. Meanwhile, RL has increasingly gained attention for its ability to adaptively learn control policies from data, allowing for more flexible and autonomous decision-making in dynamic environments [35]. Stochastic optimization approaches have also been applied to account for uncertainties in traffic behavior and communication, as seen in the work of Li [36, 37].
@@
Field experiments can further demonstrate the effectiveness of CAVs and platooning in increasing traffic stability and reducing fuel consumption. Various configurations, such as a single CAV leading multiple HDVs on a ring road [38], three trucks on a test track [39], and 100 CAVs on a freeway network [40], have been explored by researchers to highlight the benefits of integrating CAVs and platooning into traffic systems. The data generated from such field experiments can also provide valuable insights for theoretical research, as demonstrated in studies like [41] and [42]. These studies underscore the potential of advanced control techniques to further enhance the performance and adaptability of CAVs and platooning, especially in mixed autonomy settings.
@@
Despite the growing trend of investigating the stability of longitudinal interactions among multiple CAV platoons, few studies have explored their implications using ODE-based CF models or in mixed traffic scenarios involving HDVs. To address this research gap, this paper propose a foundational framework for multiple platoons of CAVs that is adaptable to various CF models and control designs. We extend a recently proposed single-platoon CF model [18] to a multi-platoon model that accommodates different control strategies and communication capabilities between platoons. Particularly, when the platoon size is set uniformly to one, the model degenerates to a classic CF model for HDVs, or AVs equipped with backward detection. Theoretically, we demonstrate that the stability of the multi-platoon models depends on the platoon size and communication level: larger platoon sizes and enhanced communication contribute to increased stability, while minimal delays between platoons have a negligible effect on stability. Numerical simulations on a ring road, involving different vehicle arrangements and mixed traffic scenarios with various sizes of CAV platoons and HDVs in different ratios and orders, support our theoretical findings. The simulation results also demonstrate that HDVs benefit from following CAV platoons, even when the CAV platoons are not specially designed to control the HDVs.
@@
The remaining part of this paper is organized as follows. In II, we introduce the CF models for single and multiple platoons of CAVs. In III, stability criteria of the proposed models are presented and proved. In section IV, we perform numerical simulations for the proposed models with various traffic assignments on a ring road. The impact of delay and connectivity are analysed. Lastly in section V, conclusion and possible extensions are given.
II Models for CAV platoons
II-A General assumptions
@@
II-B Single platoon: base model
@@
However, the reliability of (3) assumes that platoon followers can precisely acquire information from the leading vehicle with no delay, which is unrealistic for a large number of CAVs. Therefore, a multi-platoon system (where one platoon follows annother) serves as a viable formation strategy for managing long strings of vehicles, helping to mitigate the effects of communication delays and instabilities in centralized controls. In the following subsections, we provide some examples of multi-platoon models where we assume that each platoon is governed by (3) and the speed of platoon leaders are controlled as inputs. We will neglect the effects of communication delays and minor adjustments within each platoon.
II-C Multi-platoon: no inter-connection
@@
II-D Multi-platoon: two-way inter-connection
@@
$\ddot{x}_{i,1}=a$ $\left[(1+p)V\left(\frac{\Delta_{i}(t-t_{d})}{N_{i}}\right)\right.$ (6)
$\left.-pV\left(\frac{\Delta_{i-1}(t-t_{d})}{N_{i-1}}\right)-\dot{x}_{i,1}\right],$
where $\Delta_{i}(t)=x_{i+1,1}(t)-x_{i,1}(t)$ is the headway between the $i$-th and $i+1$-th platoon leader, $p$ is the smoothing factor for the platoon in the back, and $t_{d}$ is the constant communication delay between platoons. For the followers we apply the same equation (5) as in the previous model. In particular, if the platoon size $N_{m}=1$ and $t_{d}=0$ for all $m$, (6) becomes equivalent to the modified OVM with autonomous vehicles in [11].
@@
III Stability analysis
In this section, we analyse the stability of the models in section II through linear stability analysis. The steady-state (equilibrium) solution of all the aforementioned models on a ring road of length $L$ with $N_{tot}=\sum_{i=1}^{m}N_{i}$ vehicles is
@@ III Stability analysis
where $h=L/N_{tot}$ is the equilibrium headway. To analyse the effect of platoon sizes, for the stability analysis we assume that for the multi-platoon models all the platoons are of a uniform size, denoted as $N$. Then for the no-connection model proposed in Subsection II-C, the following stability criterion holds:
Theorem III.1.
@@
Using similar approaches, for the two-way connected model proposed in Subsection II-D, the following stability criterion applies:
Theorem III.2.
The two-way connected multi-platoon model (6, 5) with identical platoon size $N$ is stable if
@@
For the connected model (6, 5) we follow the assumptions of previous works, e.g. [43, 44] such that higher orders of the constant delay $t_{d}$ are neglected. And after linearization, for the connected multi-platoon system in Subsection II-D, The $m$ platoon leaders form a system of $m$ linear ODEs:
$\ddot{y}_{i,N}(t)=a$ $\left[V^{\prime}(h)(1+p)\frac{\Delta_{i,N}(t-t_{d})}{N}\right.$ (23)
$\left.-V^{\prime}(h)p\frac{\Delta_{i-1,N}(t-t_{d})}{N}-\dot{y}_{i,N}(t)\right].$
@@
We observe that the connected model exhibits larger stability regions than the non-connected model. However, the difference diminishes as platoon size increases. Figure 7 is the plot of the two-way connected model’s stability regions with different delays of backward sensitivity $p=0.3$ and platoon size $N=4$.
Figure 7: Neutral stability lines of the multi-platoon model with two-way connection of platoon size $N=4$ and delay $t_{d}=0,0.4,0.8,1.2,1.6$s.
@@
IV Numerical simulations
IV-A General information
We use MATLAB 2024a for both simulation and plots. All the simulations are performed on a single-lane ring road. To acquire more realistic results, we modified the OVM by adding a maximum acceleration constraint and an emergency braking system as follows:
@@ IV-A General information
IV-B Experiments of identical CAV platoons
@@
IV-B1 Different platoon sizes without delay
In this simulation, we aim to show the effect of platoon size and connectivity without delay. The platoon sizes are selected as $N=2,3,4,5$, and the connectivity options between platoons include no-connection, front-connection, and two-way connection. Figure 8 is the headway plots for the no-connection model with platoon size $N=2,3,4$ during $3800$s to $4000$s, and for $N=5$ from 00s to $200$s. In these plots, 20 vehicles are selected at even intervals, starting from the 1st to the 120th vehicle, with every 6th vehicle chosen for representation.
(a) $N=2$
(b) $N=3$
(c) $N=4$
(d) $N=5$
@@ IV-B1 Different platoon sizes without delay
For the headway plots in this section, if the vehicles have not stabilized after $4000$s, we select every 6th vehicle from the 1st to the 120th for plotting, with the time interval spanning from 3800 to 4000 seconds. If the vehicles have stabilized, the headway plots will instead cover the time period from 0s until stabilization (60, 200, or 300s, depending on the scenario).
@@
From the simulation results, we observe that stability of CAV platoons can be improved by increasing both platoon size and connectivity. Moreover, with two-way connections, the equilibrium state can be achieved with platoons consisting of just two CAVs. However, this condition is only guaranteed if the intra-platoon connections are robust and the inter-platoon communication is without delay.
IV-B2 Fixed platoon size with different delays
In this simulation, we aim to find the effects of inter-platoon communication delay. We fix the platoon size at $N=4$ and select different constant delays for both the front-connected model and two-way connected model. Figure 11 is the headway plots for the two-way connected model with communication delays $t_{d}=0.4,0.8,1.2,1.6$s.
(a) $t_{d}=0.4$s
(b) $t_{d}=0.8$s
(c) $t_{d}=1.2$s
(d) $t_{d}=1.6$s
Figure 11: Plots of headways for selected vehicles with two-way connection for platoon size $N=4$ and communication delays $t_{d}=0.4,0.8,1.2,1.6$s.
@@ IV-B2 Fixed platoon size with different delays
(a) $t_{d}=0.4$s
(b) $t_{d}=0.8$s
(c) $t_{d}=1.2$s
(d) $t_{d}=1.6$s
Figure 12: Plots of minimum and maximum speeds of platoons with no-connection of size $N=4$ and communication delays $t_{d}=0.4,0.8,1.2,1.6$s.
From the simulation results, we observe that increasing communication delays between platoons negatively impacts the stability of CAV platoons. Moreover, the variance in headways grows exponentially with increased delay, which aligns with theoretical analysis. These findings can guide CAV manufacturers in setting standards for sensors and other hardware components that influence communication delays.
IV-C Experiments of CAV platoons mixed with HDVs
One potential benefit of implementing CAV platoons in traffic flow is their stabilizing effect when mixed with HDVs (which can be treated as CAV platoons of size 1 with no communication). In this subsection, we test various distributions of CAV platoons and HDVs on the ring road described in Subsection IV-A to evaluate their impact on traffic flow stability.
IV-C1 Segregated CAV platoons and HDVs
@@
IV-C2 Evenly mixed CAV platoons and HDVs
@@
V Conclusions
In this paper, we have extended a recently proposed single platoon CF model to accommodate multiple platoons. By treating the leading vehicle of each platoon considered as control inputs, we developed two distinct control designs that account for varying degrees of connectivity between platoons. We showed that our proposed multi-platoon models are consistent with foundational car-following models when the platoon size is reduced to $1$.
Through linear stability analysis, we demonstrated that both platoon size and the level of inter-platoon communication can enhance system stability. The results of numerical experiments with varied platoon size and connectivity are consistent with theoretical analysis. Furthermore, when testing configurations that mixed CAV platoons with HDVs, we observed that HDVs benefit from following CAV platoons—even without specific design considerations for HDV control.
A notable outcome of our analysis is that the influence of inter-platoon connections diminishes as platoon sizes increase. This suggests that, from a manufacturing standpoint, enhancing V2V communication within platoons should be prioritized over V2I communications. Another finding is that the stability of mixed traffic flow exhibits similar characteristics in scenarios where CAV platoons are either evenly distributed or segregated—a result that consists with a study of macroscopic models of mixed flow [45].
This paper provides a solid foundation for future innovations in CAV technologies and opens several avenues for further exploration. Integration with other control strategies, such as feedback and optimal control, could significantly enhance stability, safety, and comfort for travelers. Additionally, addressing fairness within the model and considering dynamic leader switching and platoon reformation could lead to more practical and equitable applications. Extending the proposed platoon models to accommodate more complex traffic scenarios, such as multi-lane roads and signalized intersections, would broaden the models’ applicability.

### v3

# Analysis of a platooned car-following model with different inter-platoon communication levels

@@ Abstract
Despite growing interest in vehicle platooning research, the effect of communication capability between platoons is not investigated to a depth of depth. In this paper, we extend a single-platoon car-following (CF) model to multi-platoon CF models for connected and autonomous vehicles (CAVs) with different inter-platoon communication capabilities. Specifically, we consider forward and backward connection availabilities with delays between platoons. Using linear stability analysis, we discovered that for identical platoons, stability increases with platoon size and connection availabilities and decreases exponentially with large delay. With maximum acceleration and emergency braking constraints integrated into the models, we performed simulations for various cases of CAV platoons and mixed autonomy with human-driven vehicles (HDVs). The simulation results for CAV platoons are consistent with theoretical analysis. The mixed autonomy experiments demonstrate that in the ring road scenario, CAV platoons can stabilize HDVs without adaptions, and the effect of distribution is marginal. Overall, this paper provides valuable insights for designing vehicle-to-vehicle (V2V) communications and managing mixed traffic scenarios.
Keywords: CAV platoon, communication capability, linear stability, numerical simulation
1 Introduction
@@
The next stage of CF models includes multi-vehicle interaction, moving beyond the classic models considering only the immediate leader, thus modeling traffic flow more realistically.CF patterns such as multi-following [9] and backward-following [10], both based on OVM, have been increasingly explored and have inspired subsequent research on mixed traffic flow [11], [12], where CAVs are designed to better stabilize traffic. In fully autonomous traffic scenarios, various platooning designs can be integrated to enhance both efficiency and safety [13, 14, 15, 16, 17, 18]. Further improvements to CF models include the addition of delay factors [19], [20], as well as the integration of control mechanisms such as safety-prioritized control [21, 22] and feedback control [23]. These extended CF models, coupled with advanced control strategies, reflect the growing trend toward improving traffic stability, efficiency, and safety. By accounting for multi-vehicle interactions and integrating autonomous systems, these models contribute significantly to enhancing traffic performance, particularly in mixed autonomy scenarios.
@@
Control strategies for groups of CAVs, particularly Adaptive Cruise Control (ACC) and Cooperative Adaptive Cruise Control (CACC), have been extensively studied to improve vehicle platooning efficiency and safety [24, 25]. ACC primarily focuses on maintaining safe distances between vehicles by adjusting speed based on sensor data, achieving better results than typical human drivers. However, it operates in a decentralized manner without relying on vehicle-to-vehicle (V2V) communication. On the other hand, CACC utilizes V2V communication to enable more precise control and coordination among CAV platoons [26, 27, 28]. Additionally, the quality of communications plays a crucial role in the stability of CAV platoons, as shown in studies on robust communication and stability analysis [29, 30, 31, 32]. These studies highlight the advantages of CACC over ACC, particularly regarding communication and coordination within platoons.
@@
Beyond ACC and CACC, advanced control strategies such as Model Predictive Control (MPC), reinforcement learning (RL), and stochastic optimization have been explored to further enhance the performance of CAV platoons in complex traffic scenarios. MPC has been widely used due to its ability to predict future states and optimize control actions over a finite horizon [33, 34, 35, 36]. Meanwhile, RL has increasingly gained attention for its ability to adaptively learn control policies from data, allowing for more flexible and autonomous decision-making in dynamic environments [37]. Stochastic optimization approaches have also been applied to account for uncertainties in traffic behavior and communication, as seen in the work of Li [38, 39].
@@
Field experiments can further demonstrate the effectiveness of CAVs and platooning in increasing traffic stability and reducing fuel consumption. Various configurations, such as a single CAV leading multiple HDVs on a ring road [40], three trucks on a test track [41], and 100 CAVs on a freeway network [42], have been explored by researchers to highlight the benefits of integrating CAVs and platooning into traffic systems. The data generated from such field experiments can also provide valuable insights for theoretical research, as demonstrated in studies such as [43, 44]. These studies underscore the potential of advanced control techniques to further enhance the performance and adaptability of CAVs and platooning, especially in mixed autonomy settings.
While car-following and platooning models have been widely studied, most existing studies focus on a single platoon or assume fixed connectivity patterns between platoons, with limited investigation into multi-platoon stability under different inter-platoon connectivity structures. Furthermore, studies on mixed traffic flow involving CAVs and HDVs typically rely on specific control strategies applied to CAVs to improve stability, which often requires processing information from multiple surrounding vehicles. However, the inherent stabilizing effects of structured platooning without complex control mechanisms remain underinvestigated.
@@
To address these research gaps, this paper proposes a multi-platoon framework that generalizes car-following models to different inter-platoon connectivity structures. We extend a recently proposed single-platoon CF model [45] to a multi-platoon framework, incorporating varying connection structures and inter-platoon communication delays. Notably, when the platoon size is set uniformly to one, the model degenerates to a classic CF model applicable for HDVs or AVs with backward detection. Theoretically, we proved that the stability of the multi-platoon models depends on the platoon sizes and communication capabilities: larger platoon sizes and enhanced connectivity contribute to increased stability, whereas the negative impact of delays between platoons become significant when they are sufficiently large. Furthermore, we conduct numerical simulations on a ring road, testing different vehicle arrangements, including various CAV platoon sizes and mixed traffic scenarios with different CAV-to-HDV ratios and distributions. The results validate our theoretical findings, and demonstrate that HDVs benefit from following CAV platoons, even when the CAV platoons are not specially designed to control the HDVs.
@@
The remaining part of this paper is organized as follows. In section 2, we introduce the CF models for single and multiple platoons of CAVs. In section 3, the stability criteria of the proposed models are presented and proved. In section 4, we perform numerical simulations for the proposed models with various traffic assignments on a ring road. The impact of delay and connectivity are analyzed. Lastly in section 5, conclusion and possible extensions are given.
2 Models for CAV platoons
2.1 General assumptions
@@
2.2 Single platoon: base model
@@
However, the reliability of (3) is based on the assumption that platoon followers can precisely acquire information from the leading vehicle without delay, which is unrealistic for excessive amount of CAVs. Therefore, a multi-platoon system (where one platoon follows another) serves as a viable formation strategy for managing long strings of vehicles, helping to mitigate the effects of communication delays and instabilities in centralized controls. In the following subsections, we provide some examples of multi-platoon CF models where each platoon is internally robust and governed by (3). We will neglect the effects of communication delays and minor adjustments within each platoon.
2.3 Multi-platoon: no inter-connection
@@
2.4 Multi-platoon: two-way inter-connection
@@
$\ddot{x}_{i,1}=a$ $\left[(1+p)V\left(\frac{\Delta_{i-1}(t-t_{df})}{N_{i-1}}\right)-pV\left(\frac{\Delta_{i}(t-t_{db})}{N_{i+1}}\right)-\dot{x}_{i,1}\right],$ (6)
where $\Delta_{i}(t)=x_{i+1,1}(t)-x_{i,1}(t)$ is the headway between the $i$-th and $i+1$-th platoon leader, $p$ is the smoothing factor for the platoon in the back, and $t_{df},\;t_{db}$ are the forward and backward communication delays (which can be non-constant) between platoons. For the followers we apply the same equation (5) as in the previous model. In particular, if the platoon size $N_{m}=1$ and $t_{d}=0$ for all $m$, (6) becomes equivalent to the modified OVM with autonomous vehicles in [11].
@@
3 Stability analysis
In this section, we analyse the stability of the models in section 2 through linear stability analysis. The steady-state (equilibrium) solution of all the aforementioned models on a ring road of length $L$ with $N_{tot}=\sum_{i=1}^{m}N_{i}$ vehicles is
@@ 3 Stability analysis
where $h=L/N_{tot}$ is the equilibrium headway. To analyse the effect of platoon sizes, for the stability analysis we assume that for the multi-platoon models all the platoons are of a uniform size, denoted as $N$. Then for the no-connection model proposed in Subsection 2.3, the following stability criterion holds:
Theorem 3.1.
@@
Using similar approaches, for the two-way connected model proposed in Subsection 2.4, the following stability criterion applies:
Theorem 3.2.
The two-way connected multi-platoon model (6, 5) with identical platoon size $N$ and constant communication delay $t_{d}=t_{df}=t_{db}$ is stable if
@@
For the connected model (6, 5) we follow the assumptions of previous works, e.g. [46, 47] such that higher orders of the constant delay $t_{d}$ are neglected. And after linearization, for the connected multi-platoon system in Subsection 2.4, The $m$ platoon leaders form a system of $m$ linear ODEs:
$\ddot{y}_{i,N}(t)=a$ $\left[V^{\prime}(h)(1+p)\frac{\Delta_{i,N}(t-t_{d})}{N}-V^{\prime}(h)p\frac{\Delta_{i-1,N}(t-t_{d})}{N}-\dot{y}_{i,N}(t)\right].$ (23)
@@
We observe that the connected model exhibits larger stability regions than the non-connected model. However, the difference diminishes as platoon size increases. Figure 7 is the plot of stability regions of the two-way connected model with different delays of backward sensitivity $p=0.3$ and platoon size $N=4$.
Figure 7: Neutral stability lines of the multi-platoon model with two-way connection of platoon size $N=4$ and delay $t_{d}=0,0.35,0.8,1.2,1.6$s.
@@
4 Numerical simulations
4.1 General information
We use MATLAB for both simulation and plots throughout this section. The simulations are performed on a single-lane roads. To acquire more realistic results, we modified the OVM by adding a maximum acceleration constraint and an emergency braking system as follows:
@@ 4.1 General information
4.2 Experiments of platooned CAVs
@@
4.2.1 Different platoon sizes
In this simulation, we aim to show the effect of platoon size and connectivity without delay. The platoon sizes are selected as $N=2,3,4,5$, and the connectivity options between platoons include no-connection, front-connection, and two-way connection. Figure 8 is the headway plots for the no-connection model with platoon size $N=2,3,4,5$ (it is not necessary to explore $N>5$ since $N=5$ is already stabilized).
@@ 4.2.1 Different platoon sizes
For the headway plots in this section, if the vehicles have not stabilized after $4000$s, we select every 6th vehicle from the 1st to the 120th for plotting (20 vehicles in total), with the time interval spanning from 3800 to 4000 seconds. If stabilized before $4000$s, the headway plots will instead cover the time period from 00s to stabilization ($60$ to $300$s, depending on the scenarios).
@@
From the simulation results, we observe that the stability of CAV platoons improves by increasing platoon size when intra-platoon communication is robust. Moreover, with front and two-way connections, the equilibrium state can be achieved with platoons consisting of just four and two CAVs, respectively. However, this condition is satisfied only if the inter-platoon communication is delay-free. These findings highlight the critical role of intra-platoon V2V communication in maintaining stability, suggesting that prioritizing strong intra-platoon connectivity is more beneficial than relying on inter-platoon communication to achieve stability.
4.2.2 Fixed platoon size with delays
In this simulation, our objective is to find the effects of inter-platoon communication delay. We fix the platoon sizes at $N=4$ and apply constant and stochastic delays for the two-way connected model. Figure 11 is the headway plots for the two-way connected model with communication delays $t_{d}=0.35,0.8,1.2,1.6$s.
Figure 11: Plots of headways for selected vehicles with two-way connection for platoon size $N=4$ and communication delays $t_{d}=0.35,0.8,1.2,1.6$s.
@@ 4.2.2 Fixed platoon size with delays
Figure 12: Plots of minimum and maximum speeds of platoons with two-way connection of size $N=4$ and communication delays $t_{d}=0.35,0.8,1.2,1.6$s.
Figure 13 is the plots of headways and extreme speeds of all vehicles corresponding to Figure 11.
Figure 13: Plots of headways and extreme speeds of platoons with two-way connection of size $N=4$ and communication delays $t_{d}=0.6+r_{d},1+r_{d}$s.
From the simulation results, we observe that increasing communication delays between platoons negatively impacts the stability of CAV platoons. Moreover, the variance in headways grows exponentially with increased delay, which aligns with theoretical analysis. The results from random delay shows that the noises from communication have negligible to marginal impacts under certain threshold. These findings can guide CAV manufacturers in setting standards for sensors and other hardware components that influence communication delays.
4.3 Experiments of mixed autonomy
One potential benefit of implementing CAV platoons in traffic flow is their stabilizing effect when mixed with HDVs (which can be treated as CAV platoons of size 1 with no communication). In this subsection, we test various distributions of CAV platoons and HDVs on the ring road described in Subsection 4.1 to evaluate their impact on traffic flow stability.
4.3.1 Segregated CAV platoons and HDVs
@@
4.3.2 Evenly mixed CAV platoons and HDVs
@@
5 Conclusions
In this paper, we have extended a recently proposed single-platoon CF model to accommodate multiple platoons. By prioritizing the leading vehicle of each platoon, we proposed two models that account for varying degrees of connectivity between platoons. We showed that our proposed multi-platoon models are consistent with the foundational CF models when the platoon size is reduced to $1$.
Through linear stability analysis, we demonstrated that both platoon size and the level of inter-platoon communication can enhance system stability. The results of numerical experiments with varied platoon size and connectivity are consistent with theoretical analysis. Furthermore, when testing configurations that mixed CAV platoons with HDVs, we observed that HDVs benefit from following CAV platoons, even without specific design considerations for HDV control. This reveals a more intrinsic stabilizing effect of structured platooning.
A notable outcome of our analysis is that the influence of inter-platoon connections diminishes as platoon sizes increase. This occurs because larger platoons act as “self-stabilizing units”, where intra-platoon V2V control is sufficient to maintain stability, reducing the need for external coordination. This suggests that in practical deployment, prioritizing robust intra-platoon V2V communication is more effective than improving inter-platoon V2I links, especially in high-density traffic scenarios. Another finding is that the stability of mixed traffic flow exhibits similar characteristics in scenarios where CAV platoons are evenly distributed or segregated, a result that consists with a study of macroscopic models of mixed flow [48].
This paper provides a solid foundational structure for future innovations in CAV technologies and opens several avenues for further exploration. Integration with other control strategies, such as feedback and optimal control, could significantly enhance stability, safety, and comfort for travelers. Additionally, addressing fairness within the model and considering dynamic leader switching and platoon reformation could lead to more practical and equitable applications. Extending the proposed platoon models to accommodate more complex traffic scenarios, such as multi-lane roads and signalized intersections, would broaden the models’ applicability.